OpenStudy (anonymous):

summation of 5^n/e^n-3n from 1 to infinity

OpenStudy (amistre64):

\[\sum_{n=1}^{\infty}\frac{5^n}{e^n-3n}\]

OpenStudy (amistre64):

is that the problem?

OpenStudy (anonymous):

Yes determine convergence

OpenStudy (anonymous):

Do I use the geometric series test?

OpenStudy (amistre64):

and i take it this is a power series that is comprised of the sequence created by the formula... as opposed to the sequence itself right?

OpenStudy (anonymous):

Yess

OpenStudy (amistre64):

maybe an integral test could be applied ...

OpenStudy (anonymous):

could i make r=(5/e)^n?

OpenStudy (amistre64):

i dont see how you could; what are the steps you would take to determine that?

OpenStudy (anonymous):

i dk i ui guessed cause e is a constant. I'll try integral

OpenStudy (anonymous):

But it's not really a good function to integrate. what would i use to integrate that? there's so many constants

OpenStudy (amistre64):

divide it all by e^n is what i think you tried there \[\sum_{n=1}^{\infty}\frac{5^n/e^n}{e^n/e^n-3n/e^n}\] \[\sum_{n=1}^{\infty}\frac{\cfrac{5^n}{e^n}}{1-\cfrac{3n}{e^n}}\] \[\sum_{n=1}^{\infty}\frac{\left(\cfrac{5}{e}\right)^n}{1-\cfrac{3n}{e^n}}\]

OpenStudy (amistre64):

this would probably conform to \(\cfrac{a}{1-r}\), but I dont have to much experience at it to be sure

OpenStudy (amistre64):

the top is a constant that does to infinity ... so my gut says it diverges, but i cant recall it correctly to be sure

OpenStudy (anonymous):

what if i multiply by the reciproical and cancel the e^n's.....

OpenStudy (anonymous):

then i have r=5/3 which would diverge....

OpenStudy (anonymous):

but my gut says that's wrong

OpenStudy (amistre64):

3n/e^n at n=inf goes to zero leaving us with: \[\frac{5}{e}\ ^n\text{ ; at n=inf}\] so the sequence doesnt converge, therefore the series doesnt converge .. id say divergant on all accounts

OpenStudy (amistre64):

or even with the correct equation :) http://www.wolframalpha.com/input/?i=sum%285^n%2F%28e^n-3n%29%29+from+1+to+infinity

OpenStudy (anonymous):

Thanks but to prove it i need to identify bn and take the limit of both an and bn. what do I set as bn?

OpenStudy (amistre64):

the only thing i recall on that is to do partial decomp; and i dont quite see how to accomplish that on this one

OpenStudy (anonymous):

bn=5^n/e^n. n-n=1....

OpenStudy (amistre64):

i think ive used up all me wits that i can afford on this one :) good luck with it